how many moles of PCl5 can be produced from 50.0 g of Cl2

1
Expert's answer
2021-07-20T06:09:03-0400

P4+ 10Cl2 → 4PCl5

M(Cl2)= 70.90 g/mol

M(PCl5)=208.24 g/mol

"n = \frac{m}{M} \\"

n(Cl2) "= \frac{50.0}{70.90}=0.705 \;mol"

According to the reaction:

n(PCl5) "= \frac{4}{10}n(Cl_2) = \frac{4}{10} \times 0.705 = 0.282 \;mol"

m(PCl5) "= 0.282 \times 208.24 = 58.74 \;g"

Answer: 58.74 g